Sol Summary – November 30th

Sol: 3
Summary Title: Farm-to-Table
Author’s name: Matt Eby
Mission Status: In-Simulation
Sol Activity Summary: EVA excursion around Pooh’s Corner, off and running with the exercise protocol, and a small harvest from the GreenHab for dinner.
Look Ahead Plan: EVA to conduct first drone flights above ground paths of recent excursions (e.g. Marble Ritual).
Anomalies in work: EVA suit disconnected power plug.
Weather: Chilly morning, with moderating temperatures in the afternoon under brilliant cloud-free sunshine. Winds fairly light throughout the day.
Crew Physical Status: Crew in good shape, minor issues from high altitude and dry environment (minor headaches).
EVA: EVA led by the Technology Officer with support from the Crew Engineer and GreenHab Officer. 2 ¾ hour duration using Oppy and Curiosity rovers, miniature RC rover deployed at two locations to mark the mirror experiment location and to scout Pooh’s corner.
Reports to be filed: Journalist, Commander, Operations, Photos, EVA report, EVA request, GreenHab
Support Requested: N/A

Commander Report – November 30th

Sol: 3
Summary Title: Operations and EVA 3
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:

Crew 269 experienced another day of beautiful weather on Mars, chilly with light high-altitude clouds and light winds. The crew enjoyed a breakfast of blueberry pancakes and chocolate chip pancakes with maple syrup. After a morning planning meeting, HSO collected crew health stats while Technology Officer completed preparations for the afternoon EVA and Engineer and Green Hab Officer performed initial inspection and parts gathering for space suit 7 repair. At 12pm, Technology Officer, Engineer, and Green Hab Officer embarked on EVA #3 to Marble Ritual, the mirror deployment location, and along Cow Dung Rd. towards Galileo Rd. The 3-hour EVA was nominal and included data collection at multiple locations using a remote control mini-ATV with cameras and a virtual reality/augmented reality (VR/AR) headset. After the EVA, HSO, CDR, XO, and Engineer completed training and workouts with the portable fitness equipment while Green Hab Officer collected the first harvest of the mission of microgreens and radish stems and included them with a dinner of seared spam with mustard glaze over scalloped potatoes.

Operations Report – November 30th

SOL: 3

Name of person filing report: Ashley Kowalski

Non-nominal systems: (1) Wire connections to plug on EVA suit #7 (updates from yesterday’s report), (2) Microwave/vent hood light

Notes on non-nominal systems:

The crew received the EVA suit #7 repair instructions. In accordance with those instructions, the Crew Engineer with the assistance of the GreenHab Officer removed the air vent successfully. Upon inspection of the lower interior of the suit, we did not locate the tiny screw that holds the “push” button pin to the charging socket. We were able to find another screw in the EVA repair kit, and we screwed in the button pin. However, upon taking off the air vent, we realized that the actual wire connections that connect to the charging socket were disconnected. Yesterday when we reported that the EVA suit was still charging, we made this assumption based on the green light that was shining on the wall plug. We realized today that that wall plug shines green even if it is not plugged into any device, thus the wires were most likely already disconnected last night. In the RAM, we found a soldering iron, solder, and wire. We feel confident that we could reconnect these wires, but would like Mission Support’s agreement with this plan prior to proceeding. If we want to use the current plug, we would need solder wick in order to clean up the current connections before resoldering new wire connections. However, we did find a new plug in the EVA repair kit, so if we use this new plug, we would not need the solder wick. Please advise us on how to proceed. Two images are attached for your reference.

Additionally, while the Crew Engineer was on EVA this afternoon with two other crew members, the Commander investigated the light under the microwave/vent hood. She used a Philips screwdriver to remove the light cover and found there is currently no light bulb inside. Thus, she could not tell what size light bulb it takes. However, she did take a picture of the appliance information and serial number from inside the microwave door (please see attached image). Can mission support please look up the model number and determine what type of light they need to order?


Spirit rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Opportunity rover used: Yes
Hours: 105.7 (Start), 105.8 (End)
Beginning charge: 100%
Ending charge: 90%
Currently charging: No

Curiosity rover used: Yes
Hours: 214.5 (Start), 214.6 (End)
Beginning charge: 100%
Ending charge: 88%
Currently charging: No

Perseverance rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

General notes on rovers: No additional notes/comments.

Summary of Hab operations:
WATER USE: 35 gallons used today, 80 gallons used in total thus far
Water (static tank): 470 gallons
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: Yes, it was emptied around ~9:45 AM.

Summary of internet: No issues with the internet. Connection is good.

Summary of suits and radios: The two walkie talkies mentioned previous reports (one from the upstairs set and one from the downstairs set) are still not working.

Summary of GreenHab operations: Additional information can be found in the GreenHab Report.
WATER USE: ~18 gallons (~3 gallons today)
Heater: On
Supplemental light: Off
Harvest: 1.55 oz (1 oz of microgreens, 0.55 oz of radish stems)

Summary of ScienceDome operations: N/A
Dual split: No heat or AC used. All systems off.

Summary of RAM operations: Soldering irons, solder, and wire were taken from the RAM. They were not used yet. The crew simply was searching for tools and materials they would potentially need to repair the wire connections for EVA suit #7. We await feedback from Mission Support prior to completing any repairs.

Summary of any observatory issues: No observatory issues.

Summary of health and safety issues: No health and safety issues.

Questions, concerns and requests to Mission Support: This is just a general note to let you know that the hot water in the shower does NOT have a delay, it came on hot almost immediately. Also, the front air lock closed much easier today when returning from the EVA.

Journalist Report – November 30th

Crew 269 Journalist Report
Date: 30 Nov 2022
Sol: 3
Author: Barbara Braun, Health and Safety Officer

Crew 269 is settling into life on Mars! This morning the crew had fresh pancakes for breakfast, courtesy of Green Hab Officer Matt Eby. After health and safety checks, three members of the crew (Tech Officer Trevor Jahn, Crew Engineer Ashley Kowalski, and Green Hab Officer Matt Eby) prepared for a long EVA to continue collecting 3D model data. The crew was able to demonstrate an augmented reality headset at various locations in the area of the Hab and the Marble Ritual rock formation. The team also checked on the placement of the mirrors deployed on Sol 1 and mapped the area around the mirrors using “Goose,” our mini-rover. Before setting out, the EVA crew was able to consult with “tech support” via the time-delay Discord server to determine how to clean its main EVA camera, which had a small piece of debris on one of its internal mirrors.

While the EVA crew conducted its mission, the three remaining crew members (Commander Kristine Ferrone, Executive Officer Alli Taylor, and Health and Safety Officer Barbara Braun) set up the compact fitness gear in the lower deck of the Hab. The HSO gave a demonstration of the fitness gear to the Commander and XO and discussed fitness goals and protocols for the mission. The team took turns working out using the fitness gear and videos and several members of the crew rewarded themselves with quick “Navy-style” showers in the Hab!

Once back from his EVA, Green Hab Officer Eby visited the Green Hab to harvest some microgreens, which he and Crew Engineer Kowalski used to supplement a Martian dinner of seared Spam (yes, Spam!) over scalloped potatoes and peas. The crew had a dessert of rice pudding and settled in to review the 3D footage from the day and to plan the next day’s adventures.

GreenHab Report – November 30th

GreenHab Officer: Matthew Eby
Environmental control: Heater on, door closed, 20% humidity.
Average temperatures: 68-deg F at time of watering. Range prior 24 hours: 68-deg F – 85-deg F
Hours of supplemental light: 0 hours
Daily water usage for crops: Approximately 3 gallons
Daily water usage for research and/or other purposes: 0 gallons
Water in Blue Tank: Approximately 230 gallons (Based on a 250 gallon capacity, visually 90-95% full)
Time(s) of watering for crops: 4:30 PM
Changes to crops: None
Narrative: Today a small hand harvest of microgreens was carried out to provide fresh greens for crew dinner and to thin out the bed to promote further growth. In addition, the remnants of the radish bed were harvested (consisting largely of small stems), in preparation for reseeding the bed. Only limited water was applied to the pea plants as part of the plan to investigate if overwatering is the cause of the observed yellowing on the leaves.
Harvest: 1.00 oz of Microgreens, 0.55 oz of Radish stems
Support/supplies needed: None needed at this time.

EVA Report – November 30th

EVA # 3
Author: Trevor Jahn, Technology Officer
Purpose of EVA: Demonstrate Augmented Reality Programs, and collect data for future EVAs using a Remote Control Rover named Goose.
Start time: 12:10 pm
End time: 2:59 pm
Narrative: Trevor, Ashely, and Matt made their way out the airlock for their 3-hour EVA today at approximately 12:10 pm, with them they brought the crew’s remote control (RC) rover, now named Goose, an augmented reality (AR) headset, and a tote bag full of support equipment for Goose and the AR headset. The trio first conducted their AR demonstration at the crew habitat to the right of the airlock door from the outside, after their demonstration was complete, they proceeded to the corner of Cow Dung Road, and the road branching off the habitat, using crew rovers Curiosity and Opportunity, where they performed the second AR demonstration. The crew continued onto Marble Ritual in the crew rovers following completion of their work at the crossroads of Cow Dung Road and the road branching off the habitat. Curiosity and Opportunity were parked on the side of Cow Dung Road close to Marble Ritual, where the crew disembarked and made their way to the deployment location of the Mirror Sample experiment, and upon locating the Mirror Samples the crew conducted another AR demonstration before returning the crew rovers and preparing Goose for deployment and data collection. Goose was remote controlled heading south along Cow Dung Road until the crew were nearly parallel with the Mirror Samples, which were due west from their location. Goose’s data collection instruments were activated at the edge of Cow Dung Road, and the crew navigated Goose to the Mirror Sample location without difficulty. Goose made a lap around the Mirror Samples before the crew turned off data collection. The EVA was well ahead of schedule at that point, and the crew took the opportunity to take additional photos, before returning to Cow Dung Road with Goose along the same path. Data collection, using Goose, resumed from Marble Ritual, and the crew continued North along the road to see if they could reach the crossroads of Cow Dung Road and Galileo Road before 2:30 pm, when they would need to return to the crew rovers and make their way back to the crew habitat and airlock. The crew did not make it to Galileo Road before reaching their turn around time, but turned around after turning off Goose’s data collection, and driving Goose back to the crew rovers. The crew made it back into the airlock with time to spare before their EVA concluded at 3:00 pm.

Destination: Marble Ritual and Galileo Road
Coordinates (use UTM WGS 84): stop 1: (~518400, ~4250900), stop 2: (~518400, 4250800), stop 3: (~518550, ~4251000), stop 4: (~518500, ~4251000), stop 5: (~518550, ~4251000), stop 6: (~518950, ~4251500), stop 7: (~518550, ~4251000), stop 8: (~518400, ~4250900)
Participants: Matt Eby, Ashley Kowalski, Trevor Jahn
Road(s) and routes per MDRS Map: Cow Dung Rd (0110) to Marble Ritual, and almost to Gateway to Candor
Mode of travel: Rovers (Curiosity and Opportunity), Walking

Copyright © The Mars Society. All rights reserved. | Main Site